# GRPO-MA: Multi-Answer Generation in GRPO for Stable and Efficient Chain-of-Thought Training

### v1

@@ 4.2.1 Preliminaries
Fix a prompt pp. Thoughts $th_{1},\dots,th_{K}$ are sampled independently from a distribution $\pi_{\theta}(\cdot\mid p)$ conditioned on the prompt pp.
@@ 4.2.3 Analysis of the Variance Structure
Additionally, let us analyze how variance changes as $K$ increases. As $K\to\infty$, by the law of large numbers, the sample variance of the true thought values $\sigma_{\mu_{R}}^{2}=\tfrac{1}{K-1}\sum_{k=1}^{K}(\mu_{R_{k}}-\mu_{\bar{R}})^{2}$ converges to the population variance $\sigma_{\pi}^{2}=\mathrm{Var}[\mu_{R_{i}}]$, which characterizes the variability of the true thought values $\mu_{R_{i}}$ across the population of thoughts sampled from the distribution $\pi_{\theta}(\cdot\mid p)$. Next, we analyze the sum by splitting it based on the Kronecker delta, $\delta_{ik}$. The single term where $k=i$ converges to a non-zero constant, since expressions like $(\delta_{ik}-\tfrac{1}{K}-\dots)$ approach 11. Conversely, the sum of the other K1K-1 terms, where $k\neq i$, vanishes because each term is of order $O(1/K^{2})$, making their total sum $O(1/K)$. This leads to the final result where the variance converges to a limit determined only by the properties of the thought $th_{i}$ itself: $\lim_{K\to\infty}\mathrm{Var}[A(th_{i})]=\tfrac{\sigma_{R_{i}}^{2}}{M\,\sigma_{\pi}^{2}}$.
@@ 5.4 Inconsistency Analysis
We quantify the inconsistency between thoughts and answers during training. For a thought $th_{i}$ with $M$ answers, if $\operatorname{sign}(A(th_{i}))\neq\operatorname{sign}(A(ans_{i,j}))$, we mark it as inconsistent. The inconsistency rate is defined as $\text{InconsistencyRate}=\tfrac{1}{KM}\sum_{i=1}^{K}\sum_{j=1}^{M}1[A(th_{i})A(ans_{i,j})<0]$, where $\bm{1}[\cdot]$ denotes the indicator function, which equals 11 if the condition inside holds and 0 otherwise.
@@ A.2.5 Diagonality Analysis of Matrices
To examine whether the assumption of independence across thoughts (i.e., a diagonal covariance matrix) holds in practice, we conducted numerical simulations and empirically estimated the covariance structure of $\overrightarrow{V}(th)=(V_{1},\dots,V_{K})$. Specifically, we generated NN independent replications of the full $K$-dimensional estimator vector, denoted $V^{(n)}$ for $n=1,\dots,N$, and computed the empirical covariance matrix:
@@ A.3.1 Math
At test time, for each test problem from AIME2024 we generate $n=100$ independent candidate outputs (“generations”). From these 100100 generations we compute the pass@k metrics for $k\in\{10,32\}$.
@@
where the answer is represented as a single integes $d$. The format reward assigns a value of 11 if and only if the output strictly follows the required format, and 0 otherwise. The accuracy reward is +1 if the predicted answer is identical to the true answer, and 0 otherwise.
@@ A.3.2 Code
At test time, for each test problem from LiveBench we generate $n=100$ independent candidate outputs (“generations”). From these 100100 generations we compute the pass@k metrics for $k\in\{10,32\}$ as described below.
@@
The format reward assigns a value of 11 if and only if the output strictly follows the required tag structure and the content within <answer> can be parsed as a syntactically valid Python program. Otherwise the format reward is 0.
@@ A.3.3 Object Detection
where NN is the total number of samples in the test set, and $\bm{1}(\cdot)$ is the indicator function. To provide a comprehensive assessment, we report the performance across four different IoU thresholds: $\tau\in\{0.5,0.6,0.7,0.8\}$.
@@
where the bounding box is represented as a list of four integers $[d,d,d,d]$. The format reward assigns a value of 11 if and only if the output strictly follows the required format, and 0 otherwise. The accuracy reward is defined as the IoU between the predicted bounding box and the ground-truth bounding box.
@@ A.3.4 Affordance Prediction
where NN is the total number of samples in the test set, $C_{\text{pred}}^{(i)}$ is the predicted 2D coordinate $(x,y)$ for the $i$-th sample, and $M_{\text{gt}}^{(i)}$ is the corresponding ground-truth affordance mask. The notation $M_{\text{gt}}^{(i)}(C_{\text{pred}}^{(i)})$ represents the value of the mask at the predicted coordinate. $\bm{1}(\cdot)$ is the indicator function, which is 1 if the condition is true and 0 otherwise.
@@
where the final answer corresponds to a 2D coordinate $[d,d]$, with $d$ denoting an integer. The format reward assigns a value of 11 if and only if the output strictly adheres to this format; otherwise, it is set to 0. The accuracy reward evaluates the correctness of the prediction by checking whether the predicted 2D point lies within the ground-truth affordance mask (i.e., a region where the mask value equals 11). If the prediction falls inside the valid region, +1 reward is given; otherwise, it is not.
@@ A.3.5 Trajectory Prediction
The reward function consists of two complementary components: a format rewardand a accuracy reward. The format reward assigns a value of 11 if and only if the output strictly adheres to this format; otherwise, it is set to 0. To measure how well the predicted trajectory $\hat{T}$ matches the ground-truth trajectory $T^{*}$, we adopt an accuracy reward following the design in ManipVLM-R1 Song et al. (2025). Specifically, the reward is defined as
@@ A.3.6 Demand Prediction
where NN is the number of samples in the test set, $C_{\text{pred}}^{(i)}$ denotes the predicted 2D coordinate $(x,y)$ for the $i$-th sample, and $M_{\text{gt}}^{(i)}$ is the ground-truth binary mask of the demanded object. The notation $M_{\text{gt}}^{(i)}(C_{\text{pred}}^{(i)})$ indicates the mask value at the predicted location. $\bm{1}(\cdot)$ is the indicator function that equals 11 if the condition holds and 0 otherwise.
@@
where the final answer corresponds to a 2D coordinate $[d,d]$, with $d$ denoting an integer. The format reward is assigned 11 if the output strictly follows this structure, and 0 otherwise. The accuracy reward is assigned if and only if the predicted coordinate lies within the ground-truth object mask. These two rewards jointly ensure syntactically valid outputs and semantic correctness.
@@ A.3.7 OCR-based VQA
where NN is the number of questions, $M$ is the number of ground-truth answers per question, $a_{ij}$ is the $j$-th ground-truth answer for the $i$-th question $q_{i}$, and $o_{q_{i}}$ is the predicted answer. $NL(\cdot)$ denotes the normalized Levenshtein distance.
@@
The format reward is 11 if the output strictly follows this structure, and 0 otherwise. The accuracy reward corresponds to the ANLS score of the predicted answer for the current question.
@@ A.3.8 Simulator-based Visual Manipulation
The evaluation metric is Success Rate, following ManipLLM’s criterion based on the manipulated object’s displacement after the scripted sucker motion. Formally, given NN trials,
@@
The reward function during GRPO training consists of a format reward and a task reward. The format reward is 11 if the output strictly follows the required structure and 0 otherwise. The task reward is 11 if the manipulation attempt succeeds according to ManipLLM’s displacement criterion and 0 otherwise. The overall reward is defined as
@@ A.6.2 Richness of Reward Signals
where $\mathbf{1}\{\cdot\}$ denotes the indicator function, which equals 11 if the condition inside holds and 0 otherwise. Here, $T$ is the total number of time steps, $t$ indexes a specific time step, $K$ is the number of thoughts, $M$ is the number of answers per thought, and $\mathrm{AccR}^{t}_{i,j}$ denotes the accuracy reward associated with the $j$-th answer under the $i$-th thought at time step $t$. A higher NoZeroRate indicates a lower proportion of advantage collapses (where collapse means all advantage values become zero), and a higher proportion of effective gradient information contribution.

### v2

@@ 4.2.1 Preliminaries
Fix a prompt $p$. Thoughts $th_{1},\dots,th_{K}$ are sampled independently from a distribution $\pi_{\theta}(\cdot\mid p)$ conditioned on the prompt $p$.
@@ 4.2.3 Analysis of the Variance Structure
Additionally, let us analyze how variance changes as $K$ increases. As $K\to\infty$, by the law of large numbers, the sample variance of the true thought values $\sigma_{\mu_{R}}^{2}=\tfrac{1}{K-1}\sum_{k=1}^{K}(\mu_{R_{k}}-\mu_{\bar{R}})^{2}$ converges to the population variance $\sigma_{\pi}^{2}=\mathrm{Var}[\mu_{R_{i}}]$, which characterizes the variability of the true thought values $\mu_{R_{i}}$ across the population of thoughts sampled from the distribution $\pi_{\theta}(\cdot\mid p)$. Next, we analyze the sum by splitting it based on the Kronecker delta, $\delta_{ik}$. The single term where $k=i$ converges to a non-zero constant, since expressions like $(\delta_{ik}-\tfrac{1}{K}-\dots)$ approach $1$. Conversely, the sum of the other K1K-1 terms, where $k\neq i$, vanishes because each term is of order $O(1/K^{2})$, making their total sum $O(1/K)$. This leads to the final result where the variance converges to a limit determined only by the properties of the thought $th_{i}$ itself: $\lim_{K\to\infty}\mathrm{Var}[A(th_{i})]=\tfrac{\sigma_{R_{i}}^{2}}{M\,\sigma_{\pi}^{2}}$.
@@ 5.4 Inconsistency Analysis
We quantify the inconsistency between thoughts and answers during training. For a thought $th_{i}$ with $M$ answers, if $\operatorname{sign}(A(th_{i}))\neq\operatorname{sign}(A(ans_{i,j}))$, we mark it as inconsistent. The inconsistency rate is defined as $\text{InconsistencyRate}=\tfrac{1}{KM}\sum_{i=1}^{K}\sum_{j=1}^{M}1[A(th_{i})A(ans_{i,j})<0]$, where $\bm{1}[\cdot]$ denotes the indicator function, which equals $1$ if the condition inside holds and 0 otherwise.
@@ A.2.5 Diagonality Analysis of Matrices
To examine whether the assumption of independence across thoughts (i.e., a diagonal covariance matrix) holds in practice, we conducted numerical simulations and empirically estimated the covariance structure of $\overrightarrow{V}(th)=(V_{1},\dots,V_{K})$. Specifically, we generated $N$ independent replications of the full $K$-dimensional estimator vector, denoted $V^{(n)}$ for $n=1,\dots,N$, and computed the empirical covariance matrix:
@@ A.3.1 Math
At test time, for each test problem from AIME2024 we generate $n=100$ independent candidate outputs (“generations”). From these $100$ generations we compute the pass@k metrics for $k\in\{10,32\}$.
@@
where the answer is represented as a single integes $d$. The format reward assigns a value of $1$ if and only if the output strictly follows the required format, and 0 otherwise. The accuracy reward is +1 if the predicted answer is identical to the true answer, and 0 otherwise.
@@ A.3.2 Code
At test time, for each test problem from LiveBench we generate $n=100$ independent candidate outputs (“generations”). From these $100$ generations we compute the pass@k metrics for $k\in\{10,32\}$ as described below.
@@
The format reward assigns a value of $1$ if and only if the output strictly follows the required tag structure and the content within <answer> can be parsed as a syntactically valid Python program. Otherwise the format reward is 0.
@@ A.3.3 Object Detection
where $N$ is the total number of samples in the test set, and $\bm{1}(\cdot)$ is the indicator function. To provide a comprehensive assessment, we report the performance across four different IoU thresholds: $\tau\in\{0.5,0.6,0.7,0.8\}$.
@@
where the bounding box is represented as a list of four integers $[d,d,d,d]$. The format reward assigns a value of $1$ if and only if the output strictly follows the required format, and 0 otherwise. The accuracy reward is defined as the IoU between the predicted bounding box and the ground-truth bounding box.
@@ A.3.4 Affordance Prediction
where $N$ is the total number of samples in the test set, $C_{\text{pred}}^{(i)}$ is the predicted 2D coordinate $(x,y)$ for the $i$-th sample, and $M_{\text{gt}}^{(i)}$ is the corresponding ground-truth affordance mask. The notation $M_{\text{gt}}^{(i)}(C_{\text{pred}}^{(i)})$ represents the value of the mask at the predicted coordinate. $\bm{1}(\cdot)$ is the indicator function, which is 1 if the condition is true and 0 otherwise.
@@
where the final answer corresponds to a 2D coordinate $[d,d]$, with $d$ denoting an integer. The format reward assigns a value of $1$ if and only if the output strictly adheres to this format; otherwise, it is set to 0. The accuracy reward evaluates the correctness of the prediction by checking whether the predicted 2D point lies within the ground-truth affordance mask (i.e., a region where the mask value equals $1$). If the prediction falls inside the valid region, +1 reward is given; otherwise, it is not.
@@ A.3.5 Trajectory Prediction
The reward function consists of two complementary components: a format rewardand a accuracy reward. The format reward assigns a value of $1$ if and only if the output strictly adheres to this format; otherwise, it is set to 0. To measure how well the predicted trajectory $\hat{T}$ matches the ground-truth trajectory $T^{*}$, we adopt an accuracy reward following the design in ManipVLM-R1 Song et al. (2025). Specifically, the reward is defined as
@@ A.3.6 Demand Prediction
where $N$ is the number of samples in the test set, $C_{\text{pred}}^{(i)}$ denotes the predicted 2D coordinate $(x,y)$ for the $i$-th sample, and $M_{\text{gt}}^{(i)}$ is the ground-truth binary mask of the demanded object. The notation $M_{\text{gt}}^{(i)}(C_{\text{pred}}^{(i)})$ indicates the mask value at the predicted location. $\bm{1}(\cdot)$ is the indicator function that equals $1$ if the condition holds and 0 otherwise.
@@
where the final answer corresponds to a 2D coordinate $[d,d]$, with $d$ denoting an integer. The format reward is assigned $1$ if the output strictly follows this structure, and 0 otherwise. The accuracy reward is assigned if and only if the predicted coordinate lies within the ground-truth object mask. These two rewards jointly ensure syntactically valid outputs and semantic correctness.
@@ A.3.7 OCR-based VQA
where $N$ is the number of questions, $M$ is the number of ground-truth answers per question, $a_{ij}$ is the $j$-th ground-truth answer for the $i$-th question $q_{i}$, and $o_{q_{i}}$ is the predicted answer. $NL(\cdot)$ denotes the normalized Levenshtein distance.
@@
The format reward is $1$ if the output strictly follows this structure, and 0 otherwise. The accuracy reward corresponds to the ANLS score of the predicted answer for the current question.
@@ A.3.8 Simulator-based Visual Manipulation
The evaluation metric is Success Rate, following ManipLLM’s criterion based on the manipulated object’s displacement after the scripted sucker motion. Formally, given $N$ trials,
@@
The reward function during GRPO training consists of a format reward and a task reward. The format reward is $1$ if the output strictly follows the required structure and 0 otherwise. The task reward is $1$ if the manipulation attempt succeeds according to ManipLLM’s displacement criterion and 0 otherwise. The overall reward is defined as
@@ A.6.2 Richness of Reward Signals
where $\mathbf{1}\{\cdot\}$ denotes the indicator function, which equals $1$ if the condition inside holds and 0 otherwise. Here, $T$ is the total number of time steps, $t$ indexes a specific time step, $K$ is the number of thoughts, $M$ is the number of answers per thought, and $\mathrm{AccR}^{t}_{i,j}$ denotes the accuracy reward associated with the $j$-th answer under the $i$-th thought at time step $t$. A higher NoZeroRate indicates a lower proportion of advantage collapses (where collapse means all advantage values become zero), and a higher proportion of effective gradient information contribution.